OpenStudy (anonymous):

FREE MEDAL Describe how to transform the quantity of the third root of x to the fourth power, to the fifth power into an expression with a rational exponent. Make sure you respond with complete sentences.

OpenStudy (anonymous):

@Mertsj mind taking a look? :)

OpenStudy (mertsj):

\[\sqrt[3]{(x^4)^5}\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

Replace te 3 with 6, 4 with 5, and 5 with 7

OpenStudy (anonymous):

I don't know what happened when I copy and pasted the problem onto here, it mixed up the numbers haha

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

\[\sqrt[6]{(x^5)^7}\]

OpenStudy (mertsj):

Is that it?

OpenStudy (anonymous):

Yes!

OpenStudy (mertsj):

\[\sqrt[6]{x ^{35}}=x ^{\frac{35}{6}}\]

OpenStudy (anonymous):

Ok so my explanation would be to multiply 5 by 7, and then put 35 over 6?

OpenStudy (anonymous):

I figured it out, thanks!

OpenStudy (mertsj):

yw